40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
|
Message ID: 16240
Entry time: Tue Jul 6 17:40:32 2021
Reply to this: 17132
|
Author: |
Koji |
Type: |
Summary |
Category: |
General |
Subject: |
Lab cleaning |
|
|
We held the lab cleaning for the first time since the campus reopening (Attachment 1).
Now we can use some of the desks for the people to live! Thanks for the cooperation.
We relocated a lot of items into the lab.
- The entrance area was cleaned up. We believe that there is no 40m lab stuff left.
- BHD BS optics was moved to the south optics cabinet. (Attachment 2)
- DSUB feedthrough flanges were moved to the vacuum area (Attachment 3)
- Some instruments were moved into the lab.
- The Zurich instrument box
- KEPCO HV supplies
- Matsusada HV supplies
- We moved the large pile of SUPERMICROs in the lab. They are around MC2 while the PPE boxes there were moved behind the tube around MC2 area. (Attachment 4)
- We have moved PPE boxes behind the beam tube on XARM behind the SUPERMICRO computer boxes. (Attachment 7)
- ISC/WFS left over components were moved to the pile of the BHD electronics.
- Front panels (Attachment 5)
- Components in the boxes (Attachment 6)
We still want to make some more cleaning:
- Electronics workbenches
- Stray setup (cart/wagon in the lab)
- Some leftover on the desks
- Instruments scattered all over the lab
- Ewaste removal |
|
|
|
|
|
|
|
|
|
|